# Generative quantum machine learning via denoising diffusion probabilistic models

### 01 Feb 2024 v3

@@ V Applications
To showcase QuDDPM’s applications, we consider a particular realization of QuDDPM with each unitary $U_{k}^{(i)}$ and $\tilde{U}_{k}$ implemented by the fast scrambling model [49]—layers of general single-qubit rotations in between homogeneous tunable entangling layers of all-to-all ZZ rotations—and hardware efficient ansatz [50]—layers of X and Y single-qubit rotations in between layers of nearest-neighbor control-Z gates—separately (see Appendix H). While the MMD distance characterization similar to Fig. 3 is presented in Appendix M, we adopt more direct measures of performance in each application.
@@ V.1 Learning correlated noise
Figure 6: Bloch visualization of the forward (a1-a3) and backward (b1-b3,c1-c3) process. (d)(e) deviation of generated states from unit circle in $X-Z$ plane. The deviation $\braket{Y}^{2}$ for forward diffusion (red), backward training (blue), backward test (green) are plotted. The shaded area show the sample standard deviation.
@@ V.2 Learning many-body phases
As a proof-of-principle, we take the simple and well-known transverse-field Ising model (TFIM) described by the Hamitonian $H_{\rm TFIM}=-\sum_{i}Z_{i}Z_{i+1}-g\sum_{i}X_{i}.$ When $g$ increases from zero, the system undergoes a phase transition from ordered ferromagnetic phase to disordered phase, with the critical point at $g=1$. The states before diffusion is chosen from ground states of $H_{\rm TFIM}$ with $g\in[0.2,0.4)$ uniformly distributed. To test the capability of QuDDPM, we utilize the magnetization, $M=(\sum_{i}Z_{i})/n,$ to identify the phase of generated states from QuDDPM, and show its distribution in Fig. 5(b). Most generated states (blue and green) of QuDDPM lives in the ferromagentic phase, and shows a sharp contrast to the random states (orange).
@@ V.3 Learning nontrivial topology
We consider ensemble of states with a ring structure—generated by applying a unitary on a single state, e.g., |$\ket{\psi_{i}}=e^{-i\bm{x}_{i}\cdot\bm{G}}\ket{0},$ which models the scenario where one encodes the classical data $\bm{x}_{i}$ onto the quantum data $\psi_{i}$, as commonly adopted in quantum machine learning to solve classical problems [53, 54, 55, 15]. We test QuDDPM with a single qubit toy example, where the generators are chosen as Pauli-$Y$ and the rotation angles uniform in $[0,2\pi)$. In the QuDDPM training, we use the Wasserstein distance [35] (see Appendix B) to cope with the nontrivial topology. The forward noisy diffusion process on the sample data and the backward denoising process for training and testing are depicted in Fig. 6. To quantitatively evaluate the performance of QuDDPM, we evaluate the deviation by Pauli-Y expectation $\braket{Y}^{2}$ in Fig. 6(d)(e), where gradual transition between zero and Haar value of $1/3$ is observed in both forward diffusion and backward denoising.
@@ Acknowledgements.
QZ and BZ acknowledges support from NSF CAREER Award CCF-2240641 and ONR N00014-23-1-2296. XC acknowledges support from NSF CAREER Award DMS-2347760.
Appendix A Appendix on details of parameters
We list hyperparameters and performance for all generative learning tasks in Table 1 for reference, and state the targeted distribution of states to generate in the following. The major codes and data of the work can be found in Ref. [36].
Generation task $n$ $n_{A}$ $L$ $T$ $N$ Cost func. Performance
@@ Appendix A Appendix on details of parameters
Clustered state (Fig. 4a and 4b) $4$ $2$ $8$ 20 (Fig. 4b) 100 (Fig. 4a) MMD see Fig. 4a and 4b
Correlated noise (Fig. 5a) $2$ $2$ $6$ $20$ $500$ MMD
@@
Many-body phase (Fig. 5b) $4$ $2$ $12$ $30$ $100$ MMD
@@
Table 1: List of hyperparameters of QuDDPM and its performance in different generative learning tasks. To test the performance after training, we randomly sample $N_{\rm te}$ random noise states, and perform the optimized backward PQC to generate the sampled data. Data set size $N_{\rm tr}=N_{\rm te}=N$. $n$ is the number of data qubit $n$ and $n_{A}$ the ancilla qubit. $L$ is the PQC depth. $T$ is the diffusion steps. For cluster state generation, we evaluate the average fidelity with the center state in each cluster, i.e. |$\ket{0}$ for single-qubit and |$\ket{0,0}$ for two-qubit (see Appendix K).
Model $n$ $n_{A}$#\## variational parameters $N$ Cost func. Performance
@@
$100$ Error probability based Cost func.
@@
Table 2: List of hyperparameters of QuDDPM, QuDT and QuGAN for generating clustered state in Fig. 3 and Fig. 14. Data set size $N_{\rm tr}=N_{\rm te}=N$. $n$ is the number of data qubit $n$ and $n_{A}$ the ancilla qubit. In performance, the mean fidelity with the center state of the cluster |$\ket{0,0}$ is $\overline{F_{0}}=\mathbb{E}_{\ket{\psi}\in\tilde{S}}|\braket{0,0}{\psi}|^{2}$ and for true data it is $\overline{F_{0,\rm data}}=0.977\pm 0.014$ (see Appendix L).
@@
Appendix B Appendix on Wasserstein distance
For pure states, choosing the quantum trace distance (equaling infidelity) $D^{2}(\ket{\phi},\ket{\psi})=1-\lvert\braket{\phi}{\psi}\rvert^{2}$, then Kantorovich’s formulation for optimal transportation is to solve the following optimization problem
@@ Appendix B Appendix on Wasserstein distance
for $p\geq 1$, where $\Pi({\cal E}_{1},{\cal E}_{2})$ is the set of admissible transport plans (i.e., couplings) of probability distributions on $V\times V$ such that $\pi(B\times V)={\cal E}_{1}(B)$ and $\pi(V\times B)={\cal E}_{2}(B)$ for any measurable $B\subset V$; namely $\Pi({\cal E}_{1},{\cal E}_{2})$ stands for all distributions with marginals as ${\cal E}_{1}$ and ${\cal E}_{2}$. The Kantorovich problem in (8) induces a metric, known as the $p$-Wasserstein distance, on the space ${\cal P}_{p}(V)$ of probability distributions on $V$ with finite $p$-th moment. In particular, the $p$-Wasserstein distance $W_{p}({\cal E}_{1},{\cal E}_{2})=\mathsf{OPT}^{1/p}$, and it has identifiability in the sense that $W_{p}({\cal E}_{1},{\cal E}_{2})=0$ if and only if ${\cal E}_{1}={\cal E}_{2}$.
$W_{2}({\cal S}_{1},{\cal S}_{2})=\min_{P}\quad$ $\langle P,C\rangle,$
s.t. $P\bm{1}_{n}=a,$
$P^{\top}\bm{1}_{m}=b,$
$P\geq 0.$
Appendix C Appendix on related works
@@
The proposed QuDDPM represents a use case of theoretical idea of quantum information scrambling [27, 28] in the forward diffusion, and its backward denoising also connects to the measurement-induced phase transitions [60]. Here we point out that our forward diffusion circuits include an actual implementation of scrambling as part of the QuDDPM algorithm, while previous papers utilize tools from the study of quantum scrambling to understand quantum neural networks [61, 62].
@@
Below, we discuss about several related works. Ref. [63] utilizes diffusion map (DM) for unsupervised learning of topological phases and [64] proposes a diffusion $K$-means manifold clustering approach based on the diffusion distance [65]. Quantum DM algorithm has also been considered [66] for potentially quantum speed-up. However, these works are not on generative learning and do not consider any denoising process. Layer-wise training [67] also attempts to divide a training problem into sub-tasks in non-generative learning; however, the performance of such strategies is limited [68]. QuDDPM integrates the division of training task and an actual noisy diffusion process to enable provable benefit in training. After the completion of our work, we became aware of a recent paper [69], where classical diffusion on classical data is implemented. Our work focuses on quantum data, provides explicit construction of quantum diffusion and quantum denoising, loss function, training strategy and error analyses, and presents several applications.
Appendix D Appendix on future directions
@@
Finally, we point out some future directions, besides various applications of QuDDPM in learning quantum systems. Our current QuDDPM architecture requires a loss function based on fidelity estimations. For large systems, fidelity estimation can be challenging to implement. Towards efficient training in large systems, alternative loss functions can be adopted. For example, one may consider adopting another quantum circuit trained for telling the ensembles apart, such as quantum convolutional neural network [4] and other circuit architecture [70]. Such an approach will combine QuDDPM and the adversarial agent in QuGAN to resolve training problem in QuGAN. Another future direction is controlled diffusion [71]: when the ensemble has special symmetry, one can restrict the forward scrambling, the backward denoising and the random noise ensemble to that symmetry. It is also an interesting open problem how to introduce a control knob such that QuDDPM can learn multiple distributions and generate states according to an input requesting one of the distributions.
@@
Besides learning quantum errors and many-body phases, quantum sensor networks [72, 73] provide another application scenario of QuDDPM. In this scenario, one sends quantum probes to sense unitary physical process; on the return side, the receiver will collect a pure state from a distribution in the ideal case. It is an open problem how QuDDPM can be adopted to provide advantage in quantum sensing.
@@
Furthermore, the distribution can also be quantum states encoding classical data, where QuDDPM can also process classical data. Benchmarking QuDDPM and previous algorithms for classical data generative learning is an open direction.
@@ Appendix F Reproducing kernel Hilbert spaces and maximum mean discrepancy
A bivariate function $F:V\times V\to\mathbb{R}$ is a positive definite kernel if $\sum_{i,j=1}^{m}c_{i}c_{j}F(\ket{\phi}_{i},\ket{\phi}_{j})\geq 0$ for all $m\geq 1,\ket{\phi}_{1},\dots,\ket{\phi}_{m}\in V$, and $c_{1},\dots,c_{m}\in\mathbb{R}$. From the Moore-Aronszajn theorem (see e.g., [75, Theorem 7.2.4]), for every symmetric and positive-definite kernel $F$, there is a unique Hilbert space ${\cal H}:={\cal H}(F)$ of real-valued functions on $V$ such that:
@@
Appendix G SWAP test
@@
Appendix H Forward and backward circuits
@@
Appendix I Additional details on distance metrics evaluation
@@
Appendix J Details of simulation
@@
Appendix K Details of clustered state generation
@@
Appendix L Benchmarks: QuDT and QuGAN
@@
Appendix M Distance characterization in QuDDPM applications

### 16 Feb 2024 v4

@@ V Applications
To showcase QuDDPM’s applications, we consider a particular realization of QuDDPM with each unitary $U_{k}^{(i)}$ and $\tilde{U}_{k}$ implemented by the fast scrambling model [62]—layers of general single-qubit rotations in between homogeneous tunable entangling layers of all-to-all $ZZ$ rotations—and hardware efficient ansatz [63]—layers of $X$ and $Y$ single-qubit rotations in between layers of nearest-neighbor control-$Z$ gates—separately (see Appendix I). While the MMD distance characterization similar to Fig. 3 is presented in Appendix N, we adopt more direct measures of performance in each application.
@@ V.1 Learning correlated noise
Figure 6: Bloch visualization of the forward (a1)-(a3) and backward (b1)-(b3),(c1)-(c3) process. (d),(e) deviation of generated states from unit circle in $X$-$Z$ plane. The deviation $\braket{Y}^{2}$ for forward diffusion (red), backward training (blue), and backward test (green) are plotted. The shaded area shows the sample standard deviation.
@@ V.2 Learning many-body phases
As a proof of principle, we take the simple and well-known transverse-field Ising model (TFIM) described by the Hamitonian $H_{\rm TFIM}=-\sum_{i}Z_{i}Z_{i+1}-g\sum_{i}X_{i}.$ When $g$ increases from zero, the system undergoes a phase transition from an ordered ferromagnetic phase to a disordered phase, with the critical point at $g=1$. The states before diffusion are chosen from ground states of $H_{\rm TFIM}$ with $g\in[0.2,0.4)$ uniformly distributed. To test the capability of QuDDPM, we utilize the magnetization, $M=(\sum_{i}Z_{i})/n,$ to identify the phase of generated states from QuDDPM, and show its distribution in Fig. 5(b). Most generated states (blue and green) of QuDDPM lives in the ferromagentic phase, and shows a sharp contrast to the random states (orange).
@@ V.3 Learning nontrivial topology
We consider the ensemble of states with a ring structure—generated by applying a unitary on a single state, e.g., |$\ket{\psi_{i}}=e^{-i\bm{x}_{i}\cdot\bm{G}}\ket{0},$ which models the scenario where one encodes the classical data $\bm{x}_{i}$ onto the quantum data $\psi_{i}$, as commonly adopted in quantum machine learning to solve classical problems [66, 67, 68, 15]. We test QuDDPM with a single qubit toy example, where the generators are chosen as Pauli $Y$ and the rotation angles are uniform in $[0,2\pi)$. In the QuDDPM training, we use the Wasserstein distance [47] (see Appendix B) to cope with the nontrivial topology. The forward noisy diffusion process on the sample data and the backward denoising process for training and testing are depicted in Fig. 6. To quantitatively evaluate the performance of QuDDPM, we evaluate the deviation by Pauli $Y$ expectation $\braket{Y}^{2}$ in Figs. 6(d) and 6(e), where gradual transition between zero and a Haar value of $1/3$ is observed in both forward diffusion and backward denoising.
@@ Acknowledgements.
Q.Z. and B.Z. acknowledges support from NSF CAREER Awards No. CCF-2240641 and No. ONR N00014-23-1-2296. X.C. acknowledges support from NSF CAREER Award No. DMS-2347760.
Appendix A On details of parameters
We list hyperparameters and performance for all generative learning tasks in Table 1 for reference, and state the targeted distribution of states to generate in the following. The major codes and data of the work can be found in Ref. [48]. We also specify the cost function, among the two choices, maximum mean discrepancy (MMD) and Wasserstein distance.
Generation task $n$ $n_{A}$ $L$ $T$ $N$ Cost function Performance
@@ Appendix A On details of parameters
Clustered state [Figs. 4(a) and 4(b)] $4$ $2$ $8$ 20 [Fig. 4(b)] 100 [Fig. 4(a)] MMD see Figs. 4
Correlated noise [Fig. 5(a)] $2$ $2$ $6$ $20$ $500$ MMD
@@
Many-body phase [Fig. 5(b)] $4$ $2$ $12$ $30$ $100$ MMD
@@
Table 1: List of hyperparameters of quantum denoising diffusion probabilistic model (QuDDPM) and its performance in different generative learning tasks. To test the performance after training, we randomly sample $N_{\rm te}$ random noise states, and perform the optimized backward PQC to generate the sampled data. Dataset size $N_{\rm tr}=N_{\rm te}=N$. $n$ is the number of data qubit $n$ and $n_{A}$ the ancilla qubit. $L$ is the PQC depth. $T$ is the diffusion steps. For cluster state generation, we evaluate the average fidelity with the center state in each cluster, i.e. |$\ket{0}$ for single qubit and |$\ket{0,0}$ for $2$ qubit (see Appendix L).
Model $n$ $n_{A}$ No. variational parameters $N$ Cost funcion Performance
@@
$100$ Error probability based Cost function
@@
Table 2: List of hyperparameters of quantum denoising diffusion probabilistic model (QuDDPM), quantum direct transport (QuDT) and quantum generative adversarial networks (QuGANs) for generating a clustered state in Fig. 3 and Fig. 14. Dataset size $N_{\rm tr}=N_{\rm te}=N$. $n$ is the number of data qubit $n$, and $n_{A}$ is the ancilla qubit. In performance, the mean fidelity with the center state of the cluster |$\ket{0,0}$ is $\overline{F_{0}}=\mathbb{E}_{\ket{\psi}\in\tilde{S}}|\braket{0,0}{\psi}|^{2}$ and for true data it is $\overline{F_{0,\rm data}}=0.977\pm 0.014$ (see Appendix M).
@@
Appendix B On Wasserstein distance
For pure states, choosing the quantum trace distance (equaling infidelity) $D^{2}(\ket{\phi},\ket{\psi})=1-\lvert\braket{\phi}{\psi}\rvert^{2}$, then Kantorovich’s formulation for optimal transportation involves solving the following optimization problem
@@ Appendix B On Wasserstein distance
for $p\geq 1$, where $\Pi({\cal E}_{1},{\cal E}_{2})$ is the set of admissible transport plans (i.e., couplings) of probability distributions on $V\times V$ such that $\pi(B\times V)={\cal E}_{1}(B)$ and $\pi(V\times B)={\cal E}_{2}(B)$ for any measurable $B\subset V$; namely $\Pi({\cal E}_{1},{\cal E}_{2})$ stands for all distributions with marginals as ${\cal E}_{1}$ and ${\cal E}_{2}$. The Kantorovich problem in (8) induces a metric, known as the $p$-Wasserstein distance, on the space ${\cal P}_{p}(V)$ of probability distributions on $V$ with finite $p$th moment. In particular, the $p$-Wasserstein distance $W_{p}({\cal E}_{1},{\cal E}_{2})=\mathsf{OPT}^{1/p}$, and it has identifiability in the sense that $W_{p}({\cal E}_{1},{\cal E}_{2})=0$ if and only if ${\cal E}_{1}={\cal E}_{2}$. More details can be found in Appendix. G.
Appendix C On related works
@@
The proposed QuDDPM represents an application of the theoretical idea of quantum information scrambling [27, 28] in the forward diffusion, and its backward denoising also connects to the measurement-induced phase transitions [69]. Here we point out that our forward diffusion circuits include an actual implementation of scrambling as part of the QuDDPM algorithm, while previous papers utilize tools from the study of quantum scrambling to understand quantum neural networks [70, 71].
@@
Below, we discuss about several related works. Reference [72] utilizes a diffusion map (DM) for unsupervised learning of topological phases and [73] proposes a diffusion $K$-means manifold clustering approach based on the diffusion distance [74]. A quantum DM algorithm has also been considered [75] for potentially quantum speedup. However, these works are not on generative learning and do not consider any denoising process. Layerwise training [76] also attempts to divide a training problem into subtasks in nongenerative learning; however, the performance of such strategies is limited [77]. QuDDPM integrates the division of training task and an actual noisy diffusion process to enable provable benefits in training. After the completion of our work, we became aware of a recent paper [78], where hybrid quantum-classical DDPM is proposed. Our work focuses on quantum data; provides explicit construction of quantum diffusion and quantum denoising, loss function, training strategy and error analyses; and presents several applications.
Appendix D On future directions
@@
Finally, we point out some future directions, besides various applications of QuDDPM in learning quantum systems. Our current QuDDPM architecture requires a loss function based on fidelity estimations. For large systems, fidelity estimation can be challenging to implement. Toward efficient training in large systems, alternative loss functions can be adopted. For example, one may consider adopting another quantum circuit trained for telling the ensembles apart, such as a quantum convolutional neural network [4] and other circuit architecture [79]. Such an approach will combine QuDDPM and the adversarial agent in QuGAN to resolve the training problem in QuGAN. Another future direction is controlled diffusion [80]: when the ensemble has special symmetry, one can restrict the forward scrambling, the backward denoising and the random noise ensemble to that symmetry. It is also an interesting open problem of how to introduce a control knob such that QuDDPM can learn multiple distributions and generate states according to an input requesting one of the distributions.
@@
Besides learning quantum errors and many-body phases, quantum sensor networks [81, 82] provide another application scenario of QuDDPM. In this scenario, one sends quantum probes to sense a unitary physical process; on the return side, the receiver will collect a pure state from a distribution in the ideal case. It is an open problem of how QuDDPM can be adopted to provide an advantage in quantum sensing.
@@
Furthermore, the data can also be quantum states encoding classical data, where QuDDPM can also process classical data. Benchmarking QuDDPM and previous algorithms for classical data generative learning is an open direction.
@@ Appendix F Reproducing kernel Hilbert spaces and maximum mean discrepancy
A bivariate function $F:V\times V\to\mathbb{R}$ is a positive definite kernel if $\sum_{i,j=1}^{m}c_{i}c_{j}F(\ket{\phi}_{i},\ket{\phi}_{j})\geq 0$ for all $m\geq 1,\ket{\phi}_{1},\dots,\ket{\phi}_{m}\in V$, and $c_{1},\dots,c_{m}\in\mathbb{R}$. From the Moore-Aronszajn theorem (see e.g., [34, Theorem 7.2.4]), for every symmetric and positive-definite kernel $F$, there is a unique Hilbert space ${\cal H}:={\cal H}(F)$ of real-valued functions on $V$ such that:
@@
Appendix G Computation of Wasserstein distance
$W_{2}({\cal S}_{1},{\cal S}_{2})=\min_{P}\quad$ $\langle P,C\rangle,$
s.t. $P\bm{1}_{n}=a,$
$P^{\top}\bm{1}_{m}=b,$
$P\geq 0.$
Appendix H SWAP test
@@
Appendix I Forward and backward circuits
@@
Appendix J Additional details on distance metrics evaluation
@@
Appendix K Details of simulation
@@
Appendix L Details of clustered state generation
@@
Appendix M Benchmarks: QuDT and QuGAN
@@
Appendix N Distance characterization in QuDDPM applications